OpenStudy (anonymous):

Mike bungee jumped from a building with eight stories. Each story was 15 4/5 feet in height. How tall was the building?

OpenStudy (triciaal):

1 story = 15 4/5 8 stories = 8 * 1 story = 8*(15 4/5)